OpenStudy (anonymous):

if a number is divisible by 6 then it is divisible by 3

OpenStudy (muzzack):

i mean true sorry ;)

OpenStudy (dan815):

yes

OpenStudy (dan815):

as 6 = 2*3 means any number divisible by 6 is also divisible by 2 and 3

OpenStudy (goformit100):

A Welcome to Open Study. Stay Studious.

OpenStudy (muzzack):

why 6*2=12 then 12/3 = 4 6*3=18 then 18/3 = 6 6*4=24 then 24/3 = 8 and so forth

OpenStudy (doc.brown):

If any number is divisible by a number that is divisible then that number is divisible by any of the multipliers of the multipliers.

OpenStudy (doc.brown):

\[ab=c\]\[xab=xc\]\[\frac{xab}{a}=\frac{xc}{a}=xb\]or\[\frac{xab}{b}=\frac{xc}{b}=xa\]